OpenStudy (anonymous):

\[23x ^{12}\sqrt{13x ^{9}} \div 17x ^{4}\sqrt{13x ^{9}}\]

OpenStudy (mr.math):

\[{23x^{12}\sqrt{13x^9} \over 17x^4{\sqrt{13x^9}}}={23 \over 17}x^8\]

OpenStudy (anonymous):

thank you

OpenStudy (mr.math):

You're welcome.